OpenStudy (anonymous):

solve by factoring or finding square roots. 16x^2=56x

OpenStudy (anonymous):

8x(2x - 7) = 0 x = 0 or x = 3.5

OpenStudy (anonymous):

ok ?

OpenStudy (anonymous):

need any more help?

OpenStudy (anonymous):

this is best done by factoring and by writing it as 16x^2 - 56x = 0 before factoring

OpenStudy (anonymous):

x^2 + =-7x

OpenStudy (anonymous):

how did u get that?